OpenStudy (anonymous):

Which of the following represents a difference of squares binomial? A) (k - 2)2 B) (k2 + 4)(k2 + 4) C) k4 - 2 D) k4 - 16

OpenStudy (aum):

Difference of squares means each term must be a perfect square. Example: a^2 - b^2. See which choice can be rewritten as difference of two squares.

OpenStudy (anonymous):

i want to say either B or D....

OpenStudy (aum):

If B were (k2 + 4)(k2 - 4) (minus instead of plus) then it can be written as a difference of two squares. But (k2 + 4)(k2 + 4) = (k2+4)^2. But D) is k^4 - 16 = (k^2)^2 - 4^2 which is a difference of two squares.

OpenStudy (anonymous):

ohhhh okok i get it

OpenStudy (aum):

Alright.